OpenStudy (anonymous):

evaluate 81^3/4 Can someone explain how to do that?

OpenStudy (unklerhaukus):

is that \[81^{3/4}\]?

OpenStudy (anonymous):

Yes

OpenStudy (unklerhaukus):

can you express 81 as a power of some number?

OpenStudy (unklerhaukus):

81 = ? ^ ??

OpenStudy (sepeario):

This is essentially the same as \[\sqrt[4]{81^{3}}\]

OpenStudy (anonymous):

9^2

OpenStudy (unklerhaukus):

good 81 = 9^2 can you also express 9 as a power of some number?

OpenStudy (anonymous):

@Sepeario I know but idk how to solve that part.

OpenStudy (anonymous):

3^2

OpenStudy (sepeario):

@UnkleRhaukus is teaching you right now how to find the 4th root of 81. Then you will have to put this to the power of 3. It doesn't really matter what order you do it.

OpenStudy (unklerhaukus):

so you have 81 = 9^2 = (3^2)^2 can you simplify this?

OpenStudy (unklerhaukus):

remember that \[(a^n)^m=a^{n\times m}\]

OpenStudy (anonymous):

3^2x2 = 3^4 = 81

OpenStudy (unklerhaukus):

very good now you have show that \[81^{3/4} = (3^4)^{3/4}\] what happens next?

OpenStudy (anonymous):

Idk :L

OpenStudy (unklerhaukus):

\[(3^4)^{3/4}=3^{4\times3/4}=3^{\dots} \]

OpenStudy (anonymous):

3^3?

OpenStudy (unklerhaukus):

YES!

OpenStudy (anonymous):

so the final answer is 3^3 = 27?

OpenStudy (unklerhaukus):

\[\color{red}\checkmark\]

OpenStudy (anonymous):

Okay perfect not bad.

OpenStudy (anonymous):

Can I try another one on my own?

OpenStudy (anonymous):

and if you can tell me if it's right or not?

OpenStudy (unklerhaukus):

sure

OpenStudy (anonymous):

16^5/4 = 2^4^5/4 so far right?

OpenStudy (unklerhaukus):

parentheses make it clearer 16^(5/4) = (2^4)^(5/4) i'm pretty sure you mean this (which is right)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it's going to be 2^5 = 32

OpenStudy (unklerhaukus):

you got it!

OpenStudy (anonymous):

Okay thanks so much bud!

OpenStudy (unklerhaukus):

\[\ddot\smile \]

OpenStudy (alekos):

Nice one rhaukus. That's how you do it